minesweeper:

Are there any pairs of positive integers a,b so that (a+b)/2, sqrt(ab), 2ab/(a+b) and sqrt((a^2+b^2)/2) (AM, GM, HM, QM) are all integers?

SmokeyBrown:

Hi, and welcome to QuestionCove! I think I can answer the question with an example: 4 is a positive integer. So, what if a=4 and b=4? (4+4)/2 = 4, a positive integer sqrt(4*4) = 4, a positive integer 2*4*4/(4+4) = 4, a positive integer sqrt((4^2+4^2)/2)) = 4, a positive integer In fact, it seems that in general, as long as "a" and "b" are the same positive integer, any of the calculations above work. You can try the calculations with any equal pair and see for yourself. The problem does not specify that "a" and "b" must be distinct, does it? Now, if "a" and "b" were not allowed to be equal, it may be a different story. Because of the way the problem is worded, I think this solution works

minesweeper:

Oops I meant to add a and b are different

SmokeyBrown:

@minesweeper wrote:
Oops I meant to add a and b are different
Aha, that changes things a bit :) Hmmm....

SmokeyBrown:

I want to say that there aren't any pairs that fit the criteria, but I'm not sure how to prove it, and I wouldn't be completely confident in that answer without a proof... I'll try thinking about it and see if I come up with anything I do have a clarifying question: when the problem mentions AM, GM, HM, and QM, do you know what those stand for?

minesweeper:

arithmetic mean, geometric mean, harmonic mean, quadratic mean

SmokeyBrown:

@minesweeper wrote:
arithmetic mean, geometric mean, harmonic mean, quadratic mean
That's helpful! I did some research and it turns out that AM, GM, HM, and QM have certain properties and relationships between them, but I'm still not seeing a way to prove what we want... Sorry, I'm a bit stumped! I hope we can get some help from someone who's more familiar with these kind of proofs than I am

OronSH:

so of course by scaling by a suitable constant it suffices to find a b such that sqrt(ab) and sqrt(2(a^2+b^2)) are integers. Try to find a solution :)

OronSH:

use a=kx^2 and b=ky^2 @minesweeper

OronSH:

this reduces to x^4+y^4=2z^2, which should be easy

minesweeper:

bruh your bad i already did that and i couldnt find solutions i checked up to x,y=600

OronSH:

sobad you couldn't even solve an open problem i'm disappointed in you

minesweeper:

SO bad alecks bad

OronSH:

oren so bad

OronSH:

now do the 55 mohs